Problem 1. Compute the subgroups (0), (1), (2), (3), and (4) of the additive group Zs. Find the greatest common divisor of thPlease answer all parts of the question!
Answer #1

2 0 Given 25 = {0,11,213,4 375 divisors of 5 are 185 co>={o} <17= {1,213,4,03 . 227 = { 214,1,3,0) <37= { 318,4,2,0) 247= {4,@@ Ged (42,76) 76 = 4a Ci) +34 42 = 34 (1) + 8 34 = 8(4) + 2 8 = 214) +o .: Ged (42176) = 2 by Eclydion division lemna. 1 0 G